What are the measures of two complementary angles if the difference is 12?
thank you so much!

Answer by user_dude2008(1862)

x-y=12
x+y=90 ----> y=90-x


x-y=12
x-(90-x)=12
x-90+x=12
2x=12+90
2x=102
x=102/2
x=51
y=90-x=90-51=39
Angles are 51 and 39